OpenStudy (anonymous):

simplify the radical: radical 144

OpenStudy (mertsj):



myininaya (myininaya):

\[\sqrt{12 \cdot 12}=\sqrt{12^2}=?\]

OpenStudy (anonymous):

so i would be radical 12 or just plain 12

OpenStudy (mertsj):

Remember that taking the square root "undoes" squaring. So whatever you square to get 144 would be the answer to the square root of 144

OpenStudy (mertsj):

Remember that taking the square root "undoes" squaring. So whatever you square to get 144 would be the answer to the square root of 144

OpenStudy (anonymous):

my options are A. radical 12 B. 12 C. 12 radical 12

OpenStudy (mertsj):

Which one do you think, sweetheart?

OpenStudy (anonymous):

im thinking A or B im not sure

OpenStudy (turingtest):

\[\sqrt{x^2}=x\]as mertsj said, radicals 'undo' the squares

myininaya (myininaya):

And that of course if x>=0 which it is here :)

OpenStudy (mertsj):

Think like this: addition and subtraction are inverse operations. If I start with 3 add 5 and then subtract 5 I end up right back where I started with 3. Squaring and taking the square root are inverse operations. If I start with 3 and square it, I will get 9. Then if I take the square root, I will end up with 3 again.

OpenStudy (mertsj):